9:36 LTE Moodle Login IMNI Jacob 21 311 15 in 1 in (ksi) (in.in.) 0.001 The member has a cross sectional area and stress-stra need to know how to get 3.4 for largest normal stress
Answer #1

А P@ В. RA RB. aft 3ft P-RA -P-RB RB 3 RA EFx=0 - Rat RBP.P-500 200 From graph E= blu - 2x10 ksi = 0.001 Dol-0 RA*LAC + Rex Llargest normal stress RA - A Doutee 300 lb 30+2 = 32 32-38) Dinner = 30 = 3:08 psi Answel shown is incorrect. There can alsopls do appreciate it by giving it a thumbs up. Thanks